# World Model on Million-Length Video And Language With Blockwise RingAttention

### v3

@@ Abstract
Current language models fall short in understanding aspects of the world not easily described in words, and struggle with complex, long-form tasks. Video sequences offer valuable temporal information absent in language and static images, making them attractive for joint modeling with language. Such models could develop a understanding of both human textual knowledge and the physical world, enabling broader AI capabilities for assisting humans. However, learning from millions of tokens of video and language sequences poses challenges due to memory constraints, computational complexity, and limited datasets. To address these challenges, we curate a large dataset of diverse videos and books, utilize the Blockwise RingAttention technique to scalably train on long sequences, and gradually increase context size from 4K to 1M tokens. This paper makes the following contributions: (a) Largest context size neural network: We train one of the largest context size transformers on long video and language sequences, setting new benchmarks in difficult retrieval tasks and long video understanding. (b) Solutions for overcoming vision-language training challenges, including using masked sequence packing for mixing different sequence lengths, loss weighting to balance language and vision, and model-generated QA dataset for long sequence chat. (c) A highly-optimized implementation with RingAttention, Blockwise Transformers, masked sequence packing, and other key features for training on millions-length multimodal sequences. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of over 1M tokens. This work paves the way for training on massive datasets of long video and language to develop understanding of both human knowledge and the multimodal world, and broader capabilities.
@@ 1 Introduction
Figure 1: LWM can answer questions over a 1 hour YouTube video. Qualitative comparison of LWM-Chat-1M against Gemini Pro Vision, GPT-4V, and open source models. Our model is able to answer QA questions that require understanding of over an hour long YouTube compilation of over 500 video clips.
@@
Figure 2: LWM can retrieval facts across 1M context with high accuracy. Needle retrieval comparisons against Gemini Pro and GPT-4 for each respective max context length – 32K and 128K. Our model performs competitively while being able to extend to 8x longer context length. Note that in order to show fine-grained results, the x-axis is log-scale from 0-128K, and linear-scale from 128K-1M.
Current approaches on modeling the world are mostly restricted to short sequences of language sequences or short sequences of images and clips  [5, 54, 55, 44, 52]. resulting in models which lack understanding about parts of the world that are hard to represent in texts or short clips, and are unable to process complex long-form language and visual tasks. Temporal structure in video sequences provides helpful information that is missing from language or far less obvious in static images and short clips. Long language sequences encode information that short sequences cannot, crucial for various applications such as long document retrieval or coding. Long videos provide a rich context that short clips cannot grasp, showing how scenes connect, the development of events, and the cause and effect of actions within the temporal dimension of the video. This exposure to diverse long language and video scenarios also broadens the AI systems to generalize across various real-world situations. By jointly modeling both long videos and books, the model can develop an understanding of both the multimodal world and long sequences of texts and videos, leading to more advanced AI systems with a multimodal understanding, capable of assisting humans in a broader range of tasks.
To learn from video and language sequences, we need to train a model that is capable of processing more than millions of tokens per sequence and train it on a very large dataset. However, modeling millions of tokens is extremely difficult due to high memory cost, computational complexity, and lack of suitable datasets. Luckily, we have Blockwise RingAttention [39, 31], a technique for scaling up context size arbitrarily without approximations or overheads, allowing for scalable training on long sequences. We curated a large dataset of videos and languages from public book and video datasets, consisting of videos of diverse activities and long-form books. Considering the high compute cost, we grow context size from a smaller 4K context to a larger 1M context size gradually to reduce this cost, and this approach performs well in extending context effectively. Furthermore, we identify challenges associated with training on video and language: we discovered that training on a mixture of video, image, and text is crucial for optimal performance, due to images represent higher visual quality, videos offer sequential information, and text retains language understanding. To achieve this, we implemented an efficient masked sequence packing to effectively train with different sequence lengths, rather than standard sequence packing mechanism. Moreover, determining the right balance between image, video, and text training is crucial for cross modality understanding, and we suggest a ratio that proved effective. Furthermore, to address the lack of long-form chat datasets, we developed a model-generated question-answering (QA) approach by using a short-context model to generate a QA dataset from books. We found this to be crucial for chat abilities over long sequences.
@@
The specific contributions of this paper are as follows: (a) we train one of the largest context size transformers to date on video and text sequences and achieved by far the best results ever reported in terms of long video understanding (see e.g., Figure 1) and long context fact retrieval (see e.g., Figure 2). (b) We discover a branch of challenges associated with training on video and text sequences, and propose solutions for them: loss weighting to balance language and vision, masked sequence packing to effectively train with different sequence lengths, and model-generated QA dataset for long sequence chat. (c) A highly-optimized, open-source implementation with RingAttention with Blockwise Transformer, masked sequence packing and other key features for millions-length multimodal training. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of 1M tokens. Our work paves the way for training on massive datasets of long video and language, and is useful for future development of AI systems with an understanding of both human knowledge and the multimodal world, and broader capabilities.
Figure 3: This figure illustrates the multimodal training of a Large World Model. Stage 1, LLM Context Extension, focuses on expanding context size using the Books3 dataset, with context size growing from 32K to 1M. Stage 2, Vision-Language Training, focuses on training on visual and video contents of varying lengths. The pie chart details the allocation of 495B tokens across images, short and long videos, and 33B tokens of text data. The lower panel shows interactive capabilities in understanding and responding to queries about complex multimodal world.
2 Overview
@@
We train a large autoregressive transformer model with a very large context window of up to one million tokens, building upon Llama2 7B [55]. To achieve this goal, we leverage several strategies: extending the context to 1M using books (Section 3), followed by joint training on long multimodal sequences, including text-image, text-video data, and books (Section 4).
Our training stages and datasets are shown in Figure 3 and the model architecture is shown in Figure 4.
@@ 3 Stage I: Learning Long-Context Language Models
This stage aims at first developing LWM-Text and LWM-Text-Chat, a set of long-context language models learned by training on progressively increasing sequence length data with RingAttention and Blockwise Transformer, and modifying positional encoding parameters to account for longer sequence lengths (see Section 3.1). The training steps for growing context size are shown in Section 3.2. In Section 3.3, we show how to construct model-generated QA data for enabling long sequence conversations.
3.1 Extending Context
Learning long-range dependencies over sequences of millions of tokens requires (1) scalable training on such long documents, as well as a need to (2) stably extend the context of our base language.
Scalable Training on Long Documents. Training on long documents becomes prohibitively expensive due to memory constraints imposed by the quadratic complexity of computing the attention weights. In order to address these computational constraints, we use the Blockwise RingAttention [39, 31] implementation that leverages block-wise transformer with sequence parallelism to theoretically extend to an infinite context, bounded only by the number of devices available. We further fuse Blockwise RingAttention with FlashAttention [15, 48] using Pallas [4] to optimize performance compared with using XLA compiler. In general, given a large enough tokens per device, the communication cost during Blockwise Transformer and RingAttention fully overlap with computation, and does not add any extra overhead.
Progressive Training on Increasing Context Length. Although our implementation allows us to train on long documents of millions of tokens, it still remains costly since the quadratic computational complexity of attention remains, where gradient step time scales roughly linearly with context size (given a fixed number of tokens per batch). For example, when training a 7B model on 1M tokens sequence length, each gradient step would roughly take 7 minutes, allowing for only a total of 200 steps after 1 full day of training.
Therefore, we adopt a training approach inspired from [28], where our model is trained on progressively longer sequence lengths, starting from 32K tokens and ending at 1M tokens in increasing powers of two. Intuitively, this allows the model to save compute by first learning shorter-range dependencies before moving onto longer sequences. By doing this, we are able to train on orders of magnitude more tokens compared to directly training on the maximum target sequence length. The progressive training of growing context sizes is shown in Figure 3.
Positional Extrapolation for Long Contexts. For extending positional embeddings on longer contexts, we adopt a simple, scaled up version of the approach explored in [46], where the $\theta$ for RoPE [50] is scaled up with context length. We generally found this approach to be a stable method to extend positional embeddings with context lengths due to its relatively simple nature of only needing to tune a single hyperparameter. We scale up the $\theta$ for RoPE along with context window sizes – the values are shown in Table 1.
@@ 3.1 Extending Context
Figure 4: LWM is a autoregressive transformer on sequences of millions-length tokens. Each frame in the video is tokenized with VQGAN into 256 tokens. These tokens are concatenated with text tokens and fed into transformers to predict the next token autoregressively. The input and output tokens’ order reflect the varied training data formats, including image-text, text-image, video, text-video, and purely text formats. The model is essentially trained in an any-to-any manner using multiple modalities. To differentiate between image and text tokens, and for decoding, we surround video and image tokens with the special delimiters <vision> and </vision>. We also include <eof> and <eov> vision tokens to mark the end of intermediate and final frames in images and videos. For simplicity, these delimiters are not shown.

### v4

@@ Abstract
Enabling long-context understanding remains a key challenge in scaling existing sequence models – a crucial component in developing generally intelligent models that can process and operate over long temporal horizons that potentially consist of millions of tokens. In this paper, we aim to address these challenges by providing a comprehensive exploration of the full development process for producing 1M context language models and video-language models, setting new benchmarks in language retrieval and new capabilities in long video understanding. We detail our long context data curation process, progressive context extension from 4K to 1M tokens, and present an efficient open-source implementation for scalable training on long sequences. Additionally, we open-source a family of 7B parameter models capable of processing long text documents and videos exceeding 1M tokens.
@@ 1 Introduction
Figure 1: Comparison of context size in state-of-the-art LLMs. Our model and concurrent work Gemini 1.5 both achieve a 1M context size, significantly outperforming other LLMs.
@@
Figure 2: Retrieval comparisons against Gemini Pro and GPT-4. Needle retrieval comparisons against Gemini Pro and GPT-4 for each respective max context length – 32K and 128K. Our model performs competitively while being able to extend to 8x longer context length. Note that in order to show fine-grained results, the x-axis is log-scale from 0-128K, and linear-scale from 128K-1M.
Enabling long-context understanding remains a key challenge in scaling existing sequence models—a crucial step toward developing generally intelligent models that can process and operate over extended temporal horizons, potentially involving millions of tokens. Current modeling approaches are predominantly limited to processing short sequences, whether in the form of language, images, or video clips (Brown et al., 2020, Touvron et al., 2023a; b, OpenAI, 2023, Brooks et al., 2024, Team et al., 2023). As a result, these models fall short when tasked with understanding complex, long-form language and visual contexts.
However, training models to process sequences that exceed millions of tokens is a significant challenge due to the high memory and computational costs, as well as the lack of long-context data. In this work, we address these challenges by leveraging Blockwise RingAttention (Liu et al., 2024, Liu and Abbeel, 2023), a technique that scales context size without approximations or overheads, enabling efficient training on long sequences. We curate an extensive dataset of long-form videos and books from public sources, covering a wide variety of activities and narrative structures. To address the scarcity of long-form conversational datasets, we developed a model-based question-answering technique, where a short-context model generates training data from books, significantly enhancing the model’s chat capabilities over long sequences. To mitigate computational costs, we gradually extended context size from an initial 4K tokens to 1M tokens, achieving a cost-effective and scalable approach for long-context modeling.
Following this, we further train our long-context language model to incorporate visual modalities, such as image and video. Contrary to existing popular vision-language models (Liu et al., 2023a, OpenAI, 2023, Chen et al., 2023a), we opt to additionally optimize next-token prediction losses for image and video (generation) with a VQGAN (Esser et al., 2021) encoder. We encountered various challenges training on mixed modalities (video, image, text). To balance their unique characteristics - sequential information, visual detail, and linguistic content - we implement an efficient masked sequence packing strategy, as well as introduce careful loss balancing to retain short context accuracy. This approach handles varying sequence lengths more effectively than standard methods. We also optimized the ratio of image, video, and text inputs in each batch, proposing an empirically effective balance for cross-modality learning. Since our model aims to model both textual and visual projections of the world through a large context window, drawing inspiration from prior work on world models (Brooks et al., 2024, Ha and Schmidhuber, 2018), we name our work as Large World Model (LWM).
@@
Our contributions are threefold: (a) we train one of the largest context size transformers to date on long text documents and videos and achieved competitive results on long video understanding and long context fact retrieval. (b) We discover a range of challenges associated with training on long sequences and propose solutions for them: masked sequence packing to effectively train with different sequence lengths and synthetic model-generating question-answering for effective attention. (c) We provide an open-source and optimized implementation for training with millions of tokens in context, as well as a family of Llama-based 1M context models capable of processing long documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of 1M tokens.
2 Method overview
@@
We train a large autoregressive transformer model with a large context window of up to one million tokens, building upon Llama2 7B (Touvron et al., 2023b). To achieve this goal, we implement a two-stage training strategy. In Stage I (Section 3), we extend the context to 1M tokens using book-length texts. This is followed by Stage II (Section 4), where we conduct joint training on diverse long multimodal sequences, incorporating text-image data, text-video data, and book-length texts. Our model architecture is the standard autoregressive transformer design, as illustrated in Figure 3. For a comprehensive overview of our training stages and the datasets employed, please refer to Figure 4.
@@ 3 Stage I: Learning Long-Context Language Models
This stage aims at first developing LWM-Text and LWM-Text-Chat, a set of long-context language models learned by training on progressively increasing sequence length data, and modifying positional encoding parameters to account for longer sequence lengths (see Section 3.1). In Section 3.2, we show how to construct model-generated question-answering data for enabling long sequence conversations.
@@
Figure 3: Model Architecture. The LWM model is an autoregressive transformer trained on sequences of multimodal tokens. Each video frame is tokenized into 256 tokens using VQGAN, while text is processed using a Byte-Pair Encoding (BPE) tokenizer. These tokens—both image and text—are combined and input into the transformer to autoregressively predict the next token. The model can handle various input-output modalities, including text, image, video, and text-video pairs. To distinguish between images and text, special tokens <vision> and </vision> are used for image and video frames, with <eof> and <eov> marking the end of these sequences. For simplicity, delimiters are not shown in the figure.
3.1 Progressive Training Towards Long Context
Figure 4: Curated dataset and training process with progressively increasing data length and complexity. The diagram outlines a two-stage training process. Stage 1 extends text-based understanding using books datasets of increasing document lengths and token counts. Stage 2 integrates vision-language training. Pie charts display token distribution, showing that images and short-frame videos dominate visual data, while mid-length text examples lead in the text corpus.
Learning long-range dependencies over sequences of millions of tokens requires (1) memory efficient training to scale to such long sequences, as well as a need to (2) compute efficient training to extend the context of our base language model. We outline our approach to these challenges, detailing our methods for training on long sequences, designs for efficiency and stability, and experimental setup.
Training on long sequences has become prohibitively expensive due to memory constraints imposed by the quadratic complexity of attention weight computations. To address these computational limitations, we leverage recent advancements in scaling context window size, particularly Blockwise RingAttention (Liu et al., 2024). This approach theoretically allows for an infinite context, bounded only by available devices. We further enhance performance by fusing it with FlashAttention (Dao et al., 2022) using Pallas (Bradbury et al., 2018) to optimize performance compared with using XLA compiler. Notably, with enough tokens per device—already a given—the communication cost during sequence parallelism is fully overlapped by computation, resulting in no additional overhead.
For better efficiency, we adopt a training approach inspired by prior research on extending context (Jin et al., 2023a), where our model is trained on progressively longer sequence lengths, starting from 32K tokens and ending at 1M tokens in increasing powers of two. Intuitively, this allows the model to save compute by first learning shorter-range dependencies before moving onto longer sequences. For extending positional embeddings to longer contexts, we adopt a simple, scaled-up version of the approach explored in Rozière et al. (2023), where the $\theta$ parameter for RoPE (Su et al., 2024) is scaled in proportion to the context length. We found this approach to be stable for extending positional embeddings with larger context lengths due to its simplicity, requiring the tuning of only a single hyperparameter. Specifically, we scale the $\theta$ parameter for RoPE alongside increases in context window sizes – the values are shown in Table 6. The progressive training of growing context sizes is shown in Figure 4.
We initialize from LLaMA-2 7B (Touvron et al., 2023b) as base language model and progressively increase the effective context length of the model across 5 stages: 32K, 128K, 256K, 512K, and 1M. For each stage, we train on different filtered versions of the Books3 dataset from The Pile (Gao et al., 2020). Table 6 details information about each training stage, such as the number of tokens, total time, and the Books3 dataset filtering constraints. Each successive run is initialized from the prior sequence length.
3.2 Model-Generated Question-Answering For Effective Context
We construct a simple question-answering dataset to develop long-context chat capabilities. First, we split documents from the Books3 dataset into fixed chunks of 1,000 tokens, feed each chunk into our short-context language model, and prompt it to generate a question-answer pair based on the content. To create longer examples (e.g., 32K tokens), we concatenate adjacent chunks and append the relevant question-answer pairs toward the end of the sequence in a chat format. The key intuition is that the model must learn to focus on any part of the context to answer the questions, as the relevant information can appear anywhere within the sequence.
For chat fine-tuning, we train each model on a mix of the UltraChat conversation dataset (Ding et al., 2023) and our custom question-answering dataset, using approximately a 7:3 ratio. We found it crucial to pre-pack the UltraChat data to the training sequence length and keep these examples separate from our question-answering data. This separation is necessary because UltraChat data generally contains a much higher proportion of loss tokens (due to densely packed, short questions in chat), whereas our question-answering data has long questions in chat thus a significantly lower percentage of loss tokens per sequence (< 1%). This difference arises from the long documents in the given context of our question-answering data, which are not included in loss calculations. Table 7 provides further training details for each run. Notably, we do not employ progressive training for any of the chat models; instead, we initialize them from their respective pretrained models at the same context length.
Summary: Stage I progressively increase sequence lengths using our curated dataset: starting with 32K tokens and gradually scaling up to 1M tokens. Model-generated question-answering data aids in learning effective long context.
3.3 Language Evaluation Results
3.3.1 Short Context Tasks
Table 1: Performance evaluation across language tasks, comparing Llama-2 7B (4K context window) and context-expanded variants of LWM-Text (32K to 1M). The results demonstrate that increasing context length does not significantly degrade performance on tasks with shorter contexts.
LWM-Text
Task / Metric Llama-2 7B 32k 128k 256k 512k 1M
arc_challenge/acc 0.40 0.43 0.45 0.44 0.44 0.43
arc_challenge/acc_norm 0.43 0.47 0.47 0.46 0.46 0.46
hellaswag/acc 0.57 0.57 0.57 0.57 0.56 0.57
hellaswag/acc_norm 0.77 0.76 0.76 0.76 0.75 0.75
mmlu 0.39 0.4 0.41 0.41 0.36 0.35
openbookqa/acc 0.32 0.33 0.31 0.32 0.33 0.30
openbookqa/acc_norm 0.44 0.44 0.44 0.43 0.41 0.41
3.3.2 Retrieval Task: Single Information
We evaluate on the popular Needle In A Haystack task (gkamradt, 2023) – more specifically an version (ArizeAI, 2023) that finds and retrieves random numbers assigned to randomized cities from the context. Figure 2 shows that we can scale to far larger contexts compared to the current best available LLMs. Figure 11 in Appendix shows nearly perfect retrieval accuracy over the entire context of our 1M context model. Appendix C shows more single needle retrieval results for our other shorter context length models.
3.3.3 Retrieval Task: Multiple Information
We additionally examine the performance of our model on more complex variant of the needle retrieval task by mixing in multiple needles, as well as trying to retrieve a specific subset of them. Figure 5 shows multi-needle retrieval results under different settings. Our model generalizes well when retrieving a single needle from multiple needles in context, with slight degradation when asked to retrieve more than one needle. Table 2 shows multi-needle comparisons, where our model is able to perform competitively or better than GPT-4 at retrieving one needle, or slightly lower performance when retrieving more than one needle. Furthermore, our model is also able to perform well and extend to longer context lengths of up to 1M tokens and far outperforms any recent shorter context baselines applies to longer sequence lengths through positional extrapolation techniques.. However, we note that we see degradation in accuracy while increasing the difficulty of the needle retrieval task, suggesting that there is still more room to improve on the 1M context utilization of our model. We believe that our released model will provide a foundation for future work on developing longer context models, as well as encourage more challenging benchmarks that contain difficult long-range tasks that require higher levels of synthesis, rather than pure fact retrieval.
Table 2: Multi Needle in a Haystack. * denotes models after the completion of this paper.
Context Length Model $N=2,R=2$ $N=4,R=1$ $N=4,R=2$
32K Gemini Pro (02/23) 0.34 0.44 0.6
GPT-4-1106 0.97 0.95 0.9
Llama-3.1-8B-Instruct* 0.87 0.95 0.93
Qwen2.5-7B-Instruct* 1.0 1.0 0.97
Mistral-7B-Instruct-v0.3* 0.98 0.85 0.83
LWM-Text-1M (Ours) 0.84 0.97 0.84
128K Gemini Pro (02/23) - - -
GPT-4-1106 0.92 0.8 0.82
Llama-3.1-8B-Instruct* 0.98 0.91 0.87
Qwen2.5-7B-Instruct* 0.98 0.80 0.90
Mistral-7B-Instruct-v0.3* 0.85 0.75 0.68
LWM-Text-1M (Ours) 0.83 0.98 0.83
1M Gemini Pro (02/23) - - -
GPT-4-1106 - - -
Llama-3.1-8B-Instruct* 0.27 0.32 0.18
Qwen2.5-7B-Instruct* 0.0 0.0 0.0
Mistral-7B-Instruct-v0.3* 0.05 0.13 0.10
LWM-Text-1M (Ours) 0.67 0.84 0.69
3.3.4 Evaluation on LOFT
Table 3: Evaluations on some benchmarks in the LOFT dataset.
{NiceTabular}
lccc Setting: 512K Context LWM (512K) GPT-4o (128K) Claude 3 Opus (200K)
Quora $\bf 0.38$ $0.23$ $0.37$
NQ $\bf 0.37$ $0.22$ $0.37$
HotPotQA $\bf 0.72$ $0.21$ $0.32$
We further evaluate our model on a coverage of the LOFT (Lee et al., 2024) dataset collection, we provides a more natural set of benchmarks that examine capabilities for long-context models in the context of document retrieval, and RAG. The benchmark includes tasks such as duplication detection (Quora  [^1] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs), document retrieval (HotpotQA (Yang et al., 2018)), and retrieval-based question-answering (NQ). Each dataset contains a corpus of 1000s of documents, and the model is asked to retrieve a set of document ids pertaining to its specific task (Quora, HotpotQA). For RAG (NQ dataset), the model is asked to answer the question using the given context. Table 3.3.4 shows evaluations results on 512K context length against various language model baselines.
Takeaway: Long context capability enables LWM to outperform state-of-the-art text models at multiple benchmarks. This demonstrates the effectiveness of our methods for enabling long context.
4 Stage II: Extending to Long-Context Vision-Language
4.1 Architectural Modifications For Vision
We use the pretrained VQGAN (Esser et al., 2021) from aMUSEd (Patil et al., 2024) that tokenizes $256\times 256$ input images to $16\times 16$ discrete tokens. Videos are tokenized by applying the VQGAN per-frame, and concatenating the codes together. In order to distinguish between modalities when generating, as well as knowing when to switch, we introduce mechanisms to mark the end of text generation / beginning of vision generation, and vice-versa. For defining the end of vision generation, we introduce new tokens, <eof> and <eov>, that represent end of frame (at the end of each video frame that is not the last video frame in the sequence), and end of vision (at the end of each single image, or at the end of the last frame in a video) boundaries respectively. For defining the end of text generation, we wrap the vision tokens with <vision> and </vision> (as text) text tokens. The model is trained with interleaved concatenations of vision and text tokens, and predicted autoregressively (see Figure 3).